OpenStudy (anonymous):

Find all solutions 5x+2y=0 7x+3y=0

OpenStudy (anonymous):

I'm guessing you mean isolating your y?

OpenStudy (anonymous):

noppe find all possible answers for x and y

OpenStudy (anonymous):

You can't?

OpenStudy (anonymous):

All possible cordinates that works for the above

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

I only got (0,0)

OpenStudy (anonymous):

Use table of values, makes it easier.

OpenStudy (anonymous):

haha nah you should use algebra

OpenStudy (anonymous):

It ends up to the same thing anyways.

OpenStudy (anonymous):

I'll do one. 5x+2y=0 2y = -5x y = -2.5x If x is 1 then y = -2.5(1) y=-2.5 (1,-2.5) If x is 2 then y =-2.5(2) y=-5 (2,-5)

OpenStudy (anonymous):

Two coordinates for this line so you can draw it.

OpenStudy (anonymous):

Don't forget about the other line, 7x + 3y = 0

OpenStudy (anonymous):

That is so wrong

OpenStudy (anonymous):

This is a university question by the way

OpenStudy (anonymous):

Jake, I know. I was onlydoing one so he could do the other.

OpenStudy (anonymous):

oh, good catch... sorry to interrupt :)

OpenStudy (anonymous):

Haha, join in. He seems to disagree.

OpenStudy (anonymous):

your solution is so bullpellet lol

OpenStudy (anonymous):

Vanc, ALL possible coordinates... There would be many.

OpenStudy (anonymous):

Copy down the exact question.

OpenStudy (anonymous):

nah there are only 2

OpenStudy (anonymous):

it is a system

OpenStudy (anonymous):

I know.

OpenStudy (anonymous):

a cordinate that would work for both at the same time

OpenStudy (anonymous):

Now you say it x_x

OpenStudy (anonymous):

That is what a system is

OpenStudy (anonymous):

\[y _{1} = y _{2}\]

OpenStudy (anonymous):

Whew . . . what is going on here?

OpenStudy (anonymous):

Math.

OpenStudy (anonymous):

5x+2y=0 7x+3y=0 Solve the first equation for x, solve the second equation for y and substitute one into the other.

OpenStudy (anonymous):

(or use elimination, or matrices, or graphing, or . . .)